# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ C.3 Minimal Assumptions
    Reflected SDE. The Skorokhod‑reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed.
@@
    Uniform ellipticity & VMO coefficients. $\underline{\nu}|\xi|^{2}\!\leq\!\xi^{\top}a_{\pi,c}\xi\!\leq\!\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO bounds.
@@
    VMO‑smallness. On every cylinder $Q_{r}\subset\mathcal{D}$ with $r\leq R_{0}$ the VMO modulus is $\leq\eta_{0}<\bar{\eta}(d,p,q)$ of Dong and Kim, (2009).
@@
    Costate band & utility. $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ on $\mathcal{D}$ and $U(c)=c^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$.
C.4 Automatic Properties of the Model
Lemma 4 (Strong concavity of the Hamiltonian).
[…]
($-\nabla^{2}_{(\pi,c)}H(t,x,y,\lambda,\eta,\pi,c)\;\succeq\;\underline{\Lambda}I_{n+1},$
where $\underline{\Lambda}>0$ depends only on $(a,b,\gamma,\Pi_{\max},\underline{\lambda},\bar{\lambda})$.
@@ C.4 Automatic Properties of the Model
HARA implies $U^{\prime\prime}(c)\leq-\kappa_{0}<0$ on the compact $c$‑interval, while the quadratic portfolio term contributes $-x^{2}\lambda\,\sigma^{\top}\Sigma\sigma\preceq-\kappa_{1}I_{n}$ since $x\in[\underline{x},\bar{x}]$ and $\Sigma\succ 0$. Summing gives the claim. ∎
Lemma 5 (Maximal regularity for $V^{\pi,c}$).
Under Assumption 6 and the VMO‑smallness $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{*}$ lie in $W^{2,1}_{q,p}(Q_{T})$ with
$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\bigr{)}.$
Proof.
Apply Dong and Kim, (2009) to the non‑divergence PDE for $V^{\pi,c}$ using the VMO modulus bound, uniform ellipticity and bounded coefficients. The same argument holds for $V^{*}$. ∎
C.5 Foundational Lemmata
Lemma 6 (FOC‑Reduction: Coefficient Gap).
Under Assumption 6 and Lemma 4,
$E_{h}\;\leq\;L_{H}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
Proof.
The strong concavity of the Hamiltonian (Lemma 4) implies that its gradient with respect to the control, $G(u;\lambda):=\nabla_{u}H(u,\lambda)$, is a strongly monotone operator in $u$ for a fixed state $\lambda$. By the Minty–Browder theorem (Minty, , 1962; Browder, , 1965), such an operator has a Lipschitz-continuous inverse. This allows us to bound the gap in controls by the gap in costates ($E_{v}$) and the FOC residual ($\varepsilon$), which after applying the control-to-coefficient Lipschitz bound, yields the stated inequality. ∎
Lemma 7 (Source‑Gap).
Under Assumption 6,
@@ C.5 Foundational Lemmata
Compactness of the costate band and the properties of HARA utility imply that both $U$ and $(U^{\prime})^{-1}$ are Lipschitz continuous on the relevant compact intervals. The result follows from a standard mean-value argument on the first-order conditions for consumption. ∎
@@
Let $W:=V^{\pi,c}-V^{*}$. The function $W$ solves a linear parabolic PDE whose source term is determined by the coefficient and source gaps ($h$ and $f$). By the Dong–Kim gradient estimate for PDEs with VMO coefficients, a core result of maximal regularity theory, the gradient of the solution is bounded by the norm of the source term, which gives $\|\nabla W\|_{L^{q,p}}\leq C_{S}(E_{h}+E_{f})$. ∎
@@
There exists $C_{v}^{\prime\prime}$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
@@
Combining the three preceding lemmas yields the recursive inequality $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, where $\alpha=C_{S}(L_{H}+C_{f}^{\prime})$. For a sufficiently small time interval $\Delta t\leq\Delta_{\star}$, PDE theory ensures that the constant $\alpha<1$. This turns the mapping $T(E_{v})=\alpha E_{v}+\beta\varepsilon$ into a contraction. By the contraction mapping principle, this recursion converges. Iterating this argument over the $\lceil T/\Delta_{\star}\rceil$ time slabs that cover the full horizon $[0,T]$ yields a final bound for $E_{v}$ that depends only on $\varepsilon$. ∎
C.6 Final Policy‑Gap Theorem
@@
$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\kappa_{1}\Delta t+\kappa_{2}/\sqrt{M},$ with constants depending on the value function regularity established in Lemma 5.
Theorem 1 (Policy Gap Bound). Let $\pi^{*}$ be the true optimal policy and $\widehat{\pi}$ be the policy generated by the P-PGDPO algorithm using a time step $\Delta t$ and batch size $M$. Under the regularity conditions specified in Assumption 6 in the Appendix, the gap between the P-PGDPO policy and the true optimum is bounded as follows:
where:
@@
The proof follows from the triangle inequality: $\|\widehat{\pi}-\pi^{*}\|\leq L_{F}\bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{*}\|\bigr{)}\leq L_{F}(\delta_{\mathrm{BPTT}}+E_{v}).$ Substituting the bounds for $\delta_{\mathrm{BPTT}}$ from Definition 1 and for $E_{v}$ from Proposition 1 yields the final result. ∎
Appendix D Detailed Experimental Setup
This appendix provides the detailed setup for the numerical experiments presented in Section 4.
D.1 Parameter Values
@@
    • Cross-correlation $\bm{\rho}_{Y}$ ($n\times k$) components drawn uniformly from $U(-0.2,0.2)$.
@@
D.2 PG-DPO Implementation Details
@@
  • Optimizer: Adam optimizer with a learning rate of $1\times 10^{-5}$.
@@
    • Time discretization steps ($N$): $20$.
    • Initial state sampling ($\eta$): For each trajectory in a batch, the initial time $t_{0}$ is drawn uniformly from $[0,T)$, initial wealth $W_{0}$ from $U(0.1,3.0)$. The initial state variables $\mathbf{Y}_{0}=(Y_{0,1},\dots,Y_{0,k})^{\top}$ are sampled such that each component $Y_{0,i}$ is drawn independently and uniformly from a single common range $[Y_{\min},Y_{\max}]$. This range is determined by the minimum lower bound and the maximum upper bound observed across all individual factors’ typical ±$\pm 3\sigma$ intervals:
@@
D.3 Benchmark Solution and Evaluation
@@
  • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix B.2) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
@@
  • Evaluation Metrics: Performance is evaluated via the Root Mean Squared Error (RMSE) between the policies generated by (i) Baseline PG-DPO ($\bm{\pi}_{\theta}$) and the benchmark ($\bm{\pi}^{*}$), and (ii) Two-Stage PG-DPO (using Eq. (11) with BPTT-estimated costates) and the benchmark ($\bm{\pi}^{*}$). RMSEs are computed for total, myopic, and hedging components.
@@
D.4 Deep BSDE Implementation Details

### v6

@@ C.4.3 Minimal Assumptions
    The Skorokhod-reflected SDE with inward normal $\mathbf{n}\in C^{1,\alpha}$ is well posed; expanding domains converge to the non-reflected solution (Karatzas and Shreve, , 1991).
@@
    Parabolicity:
    • (U)
      Uniform ellipticity: $\underline{\nu}|\xi|^{2}\leq\xi^{\top}a_{\pi,c}\xi\leq\overline{\nu}|\xi|^{2}$ and $(a_{\pi,c},b_{\pi,c})$ have VMO/BMO (bounded mean oscillation) bounds on $\mathcal{D}$ (Dong and Kim, , 2009);
    • (H)
      The $Y$-block is uniformly elliptic with VMO bounds while the $X$-block may degenerate; MR and gradient bounds below are applied on interior cylinders $Q_{r}\Subset\mathcal{D}$ (Lieberman, , 1996).
@@
    There exists $R_{0}>0$ such that $\sup_{0<r\leq R_{0}}\omega_{a}(r)\leq\eta_{0}<\bar{\eta}(d,p,q)$ as in Dong and Kim, (2009).
@@
    Costate and curvature bands on $\mathcal{D}$: $\underline{\lambda}\leq\lambda\leq\bar{\lambda}$ and $\underline{\kappa}_{x}\leq-\partial_{x}\lambda\leq\bar{\kappa}_{x}$.
  5. 5.
    $U(c)=c^{1-\gamma}/(1-\gamma)$ with $\gamma>0$, $\gamma\neq 1$, and coefficients obey linear growth/boundedness.
Remark 1 (Ellipticity and interior estimates).
Regime (U) can be enforced by a local lower bound on $\|\pi\|$ on $\mathcal{D}$. Under (H), arguments apply on $Q_{r}\Subset\mathcal{D}$ with local constants; a finite cover yields global bounds (Lieberman, , 1996).
C.4.4 Automatic Properties of the Model
Lemma 4 (Strong monotonicity of the portfolio FOC map).
Define
$\mathcal{G}_{\pi}(\pi;c,\lambda,\partial\lambda)\;:=\;\lambda(\mu-r\mathbf{1})+\sigma\,Z_{\lambda X}^{\top}\;=\;X\bigl{(}-\partial_{x}\lambda\bigr{)}\,\Sigma\,\pi\;+\;\lambda(\mu-r\mathbf{1})\;+\;\sigma\rho\sigma_{Y}^{\top}(\partial_{Y}\lambda)^{\top}.$
If $\Sigma\succ 0$ and Assumption 5 holds, then $\partial_{\pi}\mathcal{G}_{\pi}=X(-\partial_{x}\lambda)\,\Sigma\succ 0$. Hence for all $\pi_{1},\pi_{2}$,
$\big{\langle}\mathcal{G}_{\pi}(\pi_{1})-\mathcal{G}_{\pi}(\pi_{2}),\,\pi_{1}-\pi_{2}\big{\rangle}\;\geq\;\underline{\kappa}\,\|\pi_{1}-\pi_{2}\|^{2},\qquad\underline{\kappa}:=\underline{x}\,\underline{\kappa}_{x}\,\lambda_{\min}(\Sigma)>0,$
so $\mathcal{G}_{\pi}$ is strongly monotone and has a Lipschitz inverse (Browder, , 1965; Minty, , 1962).
Lemma 5 (Maximal regularity for $V^{\pi,c}$).
Under Assumption 5 and $\eta_{0}<\bar{\eta}(d,p,q)$, both $V^{\pi,c}$ and $V^{\ast}$ lie in $W^{2,1}_{q,p}(Q_{T})$ and
$\|D^{2}V^{\pi,c}\|_{L^{q,p}}+\|\partial_{t}V^{\pi,c}\|_{L^{q,p}}\;\leq\;C_{\mathrm{MR}}\Bigl{(}\|f_{\pi,c}\|_{L^{q,p}}+\|g\|_{W^{2-2/p}_{q}}\Bigr{)},$
with the same bound for $V^{\ast}$; under (H) the estimate holds on $Q_{r}\Subset\mathcal{D}$ (Dong and Kim, , 2009).
C.4.5 Foundational Lemmas
Lemma 6 (FOC reduction: coefficient gap).
Under Assumption 5 and Lemma 4,
$E_{h}\;\leq\;L_{G}\,E_{v}\;+\;C_{\mathrm{FOC}}\,\varepsilon.$
@@ C.4.5 Foundational Lemmas
By Lemma 4, $\pi\mapsto\mathcal{G}_{\pi}(\pi;\lambda,\partial\lambda)$ has a Lipschitz inverse (Browder, , 1965; Minty, , 1962). Thus $\|\pi-\pi^{\ast}\|\leq c_{1}\bigl{(}\|\nabla V^{\pi,c}-\nabla V^{\ast}\|+\|\mathcal{R}_{\mathrm{FOC}}\|\bigr{)}$. Lipschitz dependence of coefficients on controls yields the claim after absorbing constants. ∎
Lemma 7 (Source gap).
Under Assumption 5 and HARA utility,
@@
On the costate band, $U$ and $(U^{\prime})^{-1}$ are Lipschitz on the relevant ranges; apply a mean-value argument to the consumption FOC (see PMP/HJB expositions in Yong and Zhou, , 2012). ∎
@@
Let $W:=V^{\pi,c}-V^{\ast}$. Then $W$ solves a linear parabolic PDE with source determined by $(h,f)$. Parabolic MR with VMO coefficients yields $\|\nabla W\|_{L^{q,p}}\leq C_{S}(\|h\|_{L^{q,p}}+\|f\|_{L^{q,p}})$ (Dong and Kim, , 2009). ∎
@@
There exists $C_{v}^{\prime\prime}>0$ such that $E_{v}\leq C_{v}^{\prime\prime}\,\varepsilon$.
@@
Combine the three lemmas to get $E_{v}\leq\alpha E_{v}+\beta\varepsilon$, $\alpha=C_{S}(L_{G}+C_{f}^{\prime})$. For some slab length $\tau_{\star}>0$ (set by MR constants), on any slab of length $\tau\leq\tau_{\star}$ one has $\alpha<1$, hence $E_{v}\leq\beta(1-\alpha)^{-1}\varepsilon$. Concatenate $\lceil T/\tau_{\star}\rceil$ slabs. This is an application of the contraction principle. ∎
C.4.6 Final Policy-Gap Theorem
@@
Let $\widehat{\lambda}$ estimate $\nabla V^{\pi,c}$ by BPTT. For time step $\Delta t$ and batch size $M$,
$\delta_{\mathrm{BPTT}}:=\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\;\leq\;\begin{cases}\kappa_{\mathrm{EM}}\,\Delta t^{1/2}+\kappa_{2}/\sqrt{M},\\ \kappa_{\mathrm{Mil}}\,\Delta t\;+\kappa_{2}/\sqrt{M}\quad\text{(with first-order acceleration),}\end{cases}$
in line with standard SDE error theory (Glasserman, , 2004).
Theorem 4.2 (Policy-Gap Bound). Let $\pi^{\ast}$ be optimal and $\widehat{\pi}$ the P-PGDPO policy trained with time step $\Delta t$ and batch size $M$. Under Assumption 5,
$\|\widehat{\pi}-\pi^{\ast}\|_{L^{q,p}}\;\leq\;C_{\text{tot}}\bigl{(}\varepsilon+\delta_{\mathrm{BPTT}}\bigr{)}.$
Equivalently,
$\|\widehat{\pi}-\pi^{\ast}\|_{L^{q,p}}\;\leq\;\begin{cases}C_{\text{tot}}\bigl{(}\varepsilon+\kappa_{\mathrm{EM}}\,\Delta t^{1/2}+\kappa_{2}/\sqrt{M}\bigr{)},\\[3.0pt] C_{\text{tot}}\bigl{(}\varepsilon+\kappa_{\mathrm{Mil}}\,\Delta t+\kappa_{2}/\sqrt{M}\bigr{)}.\end{cases}$
@@
By the triangle inequality,
$\|\widehat{\pi}-\pi^{\ast}\|\;\leq\;L_{F}\Bigl{(}\|\widehat{\lambda}-\nabla V^{\pi,c}\|+\|\nabla V^{\pi,c}-\nabla V^{\ast}\|\Bigr{)}\;\leq\;L_{F}\bigl{(}\delta_{\mathrm{BPTT}}+E_{v}\bigr{)},$
and substitute Definition 2 and Proposition 1. ∎
The preceding assumptions, lemmas, and theorem complete the core logical chain for the policy-gap bound. What follows is not part of the minimal existence–uniqueness–bound argument itself, but collects a number of complementary considerations and implementation notes that help ensure the assumptions are met in practice and the results can be applied robustly in computational or extended theoretical settings. These points are offered as guidance for verification, numerical implementation, and consistency across related parts of the manuscript.
  1. 1.
    VMO smallness (how to guarantee). Two routes:
    1. (a)
      Top-level assumption. State explicitly that the $x$-VMO modulus of leading coefficients satisfies $\sup_{0<r\leq R_{0}}\omega_{a}(r)\leq\eta_{0}<\bar{\eta}(d,p,q)$.
    2. (b)
      Verification from primitives. If $y\mapsto\Sigma(y)$ and $y\mapsto\sigma_{Y}(y)$ are uniformly continuous with moduli $\omega_{\Sigma}(r),\omega_{\sigma_{Y}}(r)\to 0$ and the policy map is locally Lipschitz on the costate band, then for small $r$,
      $\omega_{a}(r)\;\leq\;C\bigl{(}\omega_{\Sigma}(r)+\omega_{\sigma_{Y}}(r)\bigr{)}.$
      If $\omega_{\Sigma}(r),\omega_{\sigma_{Y}}(r)\leq C_{0}r^{\alpha}$, choose $R_{0}\leq(\eta_{0}/(CC_{0}))^{1/\alpha}$.
    See Dong and Kim, (2009) for the VMO framework in parabolic equations.
  2. 2.
    Norm coherence. All bounds are in $L^{q,p}(Q_{T})$ / $W^{2,1}_{q,p}(Q_{T})$. If elsewhere an equilibrium operator $T:X\to X^{\ast}$ is used, fix $X=W^{1,p}(\Omega)$ with
    $\|u\|_{X}\;=\;\bigl{(}\|u\|_{L^{p}}^{p}+\|\nabla u\|_{L^{p}}^{p}\bigr{)}^{1/p},$
    and state coercivity/monotonicity in this norm (avoid using only $\|u\|_{L^{p}}$). See Adams and Fournier, (2003) for Sobolev space conventions.
  3. 3.
    Operator interchanges (Fubini–Tonelli/DCT). With bounded controls and reflection on $C^{1,1}$ domains, moments are finite, so for $\varphi\in C^{1}$ of polynomial growth and integrable weight $\psi$,
    $\frac{d}{dt}\,\mathbb{E}\bigl{[}\psi(t)\,\varphi(t,X_{t},Y_{t})\bigr{]}\;=\;\mathbb{E}\bigl{[}\psi^{\prime}(t)\,\varphi(t,X_{t},Y_{t})\bigr{]}\;+\;\mathbb{E}\bigl{[}\psi(t)\,(\partial_{t}+\mathcal{L})\,\varphi(t,X_{t},Y_{t})\bigr{]},$
    and dominated convergence justifies interchanging $\int_{0}^{T}$ and $\mathbb{E}[\cdot]$. See Karatzas and Shreve, (1991) for reflected SDE background.
  4. 4.
    Numerical convergence (strong vs. weak). For Euler–Maruyama with step $h$ and smooth coefficients:
    $\bigl{(}\mathbb{E}\|Z_{T}^{\mathrm{EM}}-Z_{T}\|^{2}\bigr{)}^{1/2}=\mathcal{O}(h^{1/2})\quad\text{(strong)},\qquad\bigl{|}\mathbb{E}\phi(Z_{T}^{\mathrm{EM}})-\mathbb{E}\phi(Z_{T})\bigr{|}=\mathcal{O}(h)\quad\text{(weak)}.$
    This matches Definition 2; Milstein/RE gives first-order in the discretization term. See Platen and Bruti-Liberati, (2010) for SDE numerical analysis.
  5. 5.
    Time-slab vs. time-step. The contraction uses an analytic slab length $\tau\leq\tau_{\star}$ tied to PDE constants; it is independent of the numerical step $\Delta t$.
  6. 6.
    Optional GE module (if used elsewhere). If you introduce $T:X\to X^{\ast}$ with
    $\langle T(u),v\rangle=\int_{\Omega}\!\Big{(}A\nabla u\cdot\nabla v+Buv+N(u)\,v\Big{)}\,dx,$
    $A\succeq\underline{\lambda}I$, $B\geq 0$, and $N$ monotone with $N(u)u\geq c_{0}|u|^{p}-C_{0}$, then
    $\langle T(u),u\rangle\;\geq\;c\,\|u\|_{W^{1,p}}^{p}-C,$
    so coercivity condition holds in the same $W^{1,p}$ norm. See Browder, (1965) for monotone operator theory.
Appendix D Classification of Parabolic HJB PDEs in Financial Models
Table 6 summarizes the classification of parabolic HJB PDEs in financial markets. Although the theoretical proofs for classes (H) and (D) are challenging, our P-PGDPO is applicable to the broader classes of parabolic PDEs from the financial and economic models.
Table 6: Classification of Parabolic HJB PDEs in Financial Models
($\implies$ Value Function / Costate very smooth) Examples: Merton’s portfolio problem with constant coefficients ($\mu,\sigma>0,r$ are constants); models with state-dependent volatility that remains strictly positive; the multi-asset OU model with one factor in Section 3.2.1 (if asset volatilities $\sigma_{i}>0$, factor volatility $\sigma_{Y}>0$, and $\Psi$ is positive definite).
($\implies$ Value Function / Costate smooth in relevant regions) Examples: The multi-asset and multi-factor model in Section 3.2.1 (if some factor volatilities in $\bm{\sigma}_{Y}$ can be zero at certain states but the overall system structure, possibly involving $\kappa_{Y}$, ensures hypo-ellipticity); a Vasicek interest rate model affecting asset returns $\mu$ but not $\sigma$.
Appendix E Detailed Experimental Setup
This section provides a detailed setup for the numerical experiments presented in Section 5. The implementation code and detailed explanations can be found at https://github.com/huhjeonggyu/PGDPO.
E.1 Parameter Values
@@
    • Cross-correlation $\bm{\rho}_{Y}$ ($n\times k$) components drawn uniformly from U(0.2,0.2)U(-0.2,0.2).
@@
E.2 PG-DPO Algorithm Details
Each iteration $j$ of the algorithm involves the following steps:
  1. Step 1.
  2. Step 2.
  3. Step 3.
  4. Step 4.
  5. Step 5.
  6. Step 6.
    where $\alpha_{j}$ is the learning rate.
By iterating steps from 1 to 6, the baseline PG-DPO algorithm gradually optimizes the policy $(\bm{\pi}_{\theta},C_{\phi})$ to maximize the extended value function $\widetilde{J}$.
E.3 PG-DPO Implementation Details
@@
  • Optimizer: Adam optimizer with a learning rate of $1.0\times 10^{-5}$.
@@
    • Time discretization steps ($N$): 2020.
    • Initial state sampling ($\eta$): For each trajectory in a batch, the initial time $t_{0}$ is drawn uniformly from $[0,T)$, initial wealth $W_{0}$ from $U(0.1,3.0)$. The initial state variables $\mathbf{Y}_{0}=(Y_{0,1},\dots,Y_{0,k})^{\top}$ are sampled such that each component $Y_{0,i}$ is drawn independently and uniformly from a single common range $Y_{\min},Y_{\max}]$. This range is determined by the minimum lower bound and the maximum upper bound observed across all individual factors’ typical $\pm 3\sigma$ intervals:
@@
E.4 Benchmark Solution and Evaluation
@@
  • Benchmark Computation: The system of ODEs (including a matrix Riccati equation) associated with the HJB equation for this terminal wealth problem (Appendix LABEL:app:analytic_MF) is solved numerically backwards from the fixed horizon $T$ using ‘scipy.integrate.solve_ivp’ (‘Radau’ method) to obtain the coefficients determining the value function and the optimal policy $\bm{\pi}^{*}(t,W_{t},\mathbf{Y}_{t})$.
@@
  • Evaluation Metrics: Performance is evaluated via the Root Mean Squared Error (RMSE) between the policies generated by (i) Baseline PG-DPO ($\bm{\pi}_{\theta}$) and the benchmark ($\bm{\pi}^{*}$), and (ii) Two-Stage PG-DPO (using Eq. (3.2) with BPTT-estimated costates) and the benchmark ($\bm{\pi}^{*}$). RMSEs are computed for total, myopic, and hedging components.
@@
E.5 Deep BSDE Implementation Details